# Function spaces for decoupling

### v2

@@ 6.3. Fractional integration
Although Theorem 6.4 only yields a strict improvement of (6.11) for $p\leq 2\leq r$, outside of this range we can nonetheless improve a classical result about the mapping properties of Fourier integral operators. Recall the definition of Fourier integral operators in standard form, from Definition 2.12.
@@
(6.13) $T:\mathcal{L}_{W,s+m+\frac{n+1}{2}(\frac{1}{p}-\frac{1}{q})}^{q,p}(\mathbb{R}^{n})\to\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$
@@
    $2\leq p\leq q<\infty$ and $m=\frac{1}{q}-\frac{1}{p}-2s(q)$.
Then $T:\mathcal{H}^{p}(\mathbb{R}^{n})\to L^{q}(\mathbb{R}^{n})$ is bounded.
@@
The first statement follows by combining Theorem 6.4 with the boundedness of $T$ from $\mathcal{L}_{W,s+m}^{q,q}(\mathbb{R}^{n})=\mathcal{H}^{s+m,q}_{FIO}(\mathbb{R}^{n})$ to $\mathcal{L}_{W,s}^{q,q}(\mathbb{R}^{n})=\mathcal{H}^{s,q}_{FIO}(\mathbb{R}^{n})$. For $m=s=0$, this invariance is contained in [19, Theorem 6.10] in cases (2) and (3), and the techniques used there also allow one to treat operators as in (1). For general $m,s\in\mathbb{R}$, the mapping property $T:\mathcal{H}^{s+m,q}_{FIO}(\mathbb{R}^{n})\to\mathcal{H}^{s,q}_{FIO}(\mathbb{R}^{n})$ can be found in [27, Proposition 3.3 and Corollary 3.4] in cases (1) and (3). In case (2) it follows from [29, Proposition 3.3], after precomposing with the operator $\langle D\rangle^{-m}$.
The second statement follows from the first statement, combined with Theorem 6.3. Indeed, in both cases one has
$\mathcal{H}^{p}(\mathbb{R}^{n})\subseteq\mathcal{L}_{W,-s(p)}^{q,p}(\mathbb{R}^{n})=\mathcal{L}_{W,s(q)+m+\frac{n+1}{2}(\frac{1}{p}-\frac{1}{q})}^{q,p}(\mathbb{R}^{n})$
so that the embedding $\mathcal{L}_{W,s(q)}^{q,p}(\mathbb{R}^{n})\subseteq L^{q}(\mathbb{R}^{n})$ concludes the proof. ∎
@@
In the case of an operator as in (1) with symbol in $S^{m}_{1,0}$, the final statement in Corollary 6.6 is well known (see e.g. [37, Section IX.6.15]), and from this one can derive the same statement for operators as in (3). Corollary 6.6 improves upon this result in several ways, since it allows for the larger class of $S^{m}_{1/2,1/2,1}$ symbols, removes the assumption that the symbol has compact spatial support, cf. (2), and yields stronger estimates through (6.13), due to the sharpness of the Sobolev embeddings in Theorem 6.3.
@@ 7.1. A discrete description of the $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$ norm
Throughout this section, for each $R\geq 2$, fix a maximal collection $V_{R}\subseteq S^{n-1}$ of unit vectors such that $|\nu-\nu^{\prime}|\geq R^{-1/2}$ for all $\nu,\nu^{\prime}\in V_{R}$. Then $V_{R}$ has approximately $R^{(n-1)/2}$ elements. Let $(\chi_{\nu})_{\nu\in V_{R}}\subseteq C^{\infty}(\mathbb{R}^{n}\setminus\{0\})$ be an associated partition of unity. More precisely, each $\chi_{\nu}$ is positively homogeneous of degree 00 and satisfies $0\leq\chi_{\nu}\leq 1$ and
@@
for all $\nu\in V_{R}$. Such a collection is straightforward to construct (see e.g. [37, Section IX.4]), and the collection $\{\mathcal{F}^{-1}(\chi_{\nu})\mid R\geq 2,\nu\in V_{R}\}$ is uniformly bounded in $L^{1}(\mathbb{R}^{n})$.
@@
The proof of (7.1) where $p=q$ is covered by [29, Proposition 4.1], at least for $R=2^{k}$ with $k\in{\mathbb{N}}$. For general $p$ and $q$, the statement follows from [30, Proposition 2.4], since $\|f\|_{L^{p}(\mathbb{R}^{n})}$ and the Besov space norm $\|f\|_{B^{0}_{p,q}(\mathbb{R}^{n})}$ are comparable, which in turn follows from the assumption that $f$ has frequency support in a dyadic annulus.
@@ 7.2. Decoupling for the sphere
For $\kappa\geq 1$ and $R\geq 2\kappa$, let $S_{R,\kappa}$ be a $\kappa R^{-1}$-neighborhood of the sphere $S^{n-1}$ in $\mathbb{R}^{n}$, and let $V_{R}$ be as in the previous subsection. Then $\theta_{\nu}:=\mathrm{supp}(\chi_{\nu})\cap S_{R,\kappa}$, for each $\nu\in V_{R}$, is a curved rectangle of dimensions approximately $R^{-1/2}\times\dots\times R^{-1/2}\times R^{-1}$ pointing in the direction of $\nu$. Moreover, $S_{R,\kappa}=\cup_{\nu\in V_{R}}\theta_{\nu}$, and the $\theta_{\nu}$ have finite overlap.
@@ 8.1. Local smoothing
For $p=q$, the statement is contained in [29, Theorem 4.4]. For $q=2$, it is a consequence of [30, Theorem 1.1] and the standard embedding $W^{s+\alpha(p)+\varepsilon,p}(\mathbb{R}^{n})\subseteq B^{s+\alpha(p)+\varepsilon/2}_{p,p}(\mathbb{R}^{n})$. For general $q\geq 2$, it follows by combining the latter result with the embedding $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})\subseteq\mathcal{L}_{W,s}^{2,p}(\mathbb{R}^{n})$ from (6.1). ∎
@@
for an implicit constant independent of $f\in W^{s+\sigma(p)+\varepsilon,p}(\mathbb{R}^{n})$. Here $\sigma(p):=0$ for $2<p\leq 2n/(n-1)$, and $\sigma(p):=2s(p)-1/p$ for $p\geq 2n/(n-1)$. By the Sobolev embeddings for $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$ in Theorem 6.3 (1), (8.1) improves upon (8.2) for $p\geq 2(n+1)/(n-1)$ and $2\leq q\leq p$. Moreover, by the sharpness of the Sobolev embedding in Theorem 6.3 (2), (8.1) in fact yields a strict improvement of (8.2) for such $p$ and $q$. On the other hand, for $2<p<2(n+1)/(n-1)$ and $2\leq q\leq p$, (8.1) neither follows from (8.2), nor vice versa. In particular, due to the sharpness of the Sobolev embedding in Theorem 6.3 (2), (8.1) yields sharper estimates than (8.2) for certain initial data. Here it is relevant to note that the exponents $\alpha(p)$ in (8.1) and $\sigma(p)$ in (8.2) are sharp, for all $2<p<\infty$ and $2\leq q\leq p$. In the case of (8.1), this follows from (6.1) and [29, Theorem 5.3].
@@
By Theorem 6.4, one has $W^{1/2,2}(\mathbb{R}^{n})\subsetneq\mathcal{L}_{W,0}^{2,p}(\mathbb{R}^{n})$ for $p=2(n+1)/(n-1)$. Given that $\alpha(p)=0$, Theorem 8.1 therefore almost yields a strict improvement of the classical Strichartz estimate
@@ 8.2. Nonlinear wave equations
In this subsection we indicate how our results can be applied to nonlinear wave equations with initial data in $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$ for $p>2$. Such initial data might be referred to as “slowly decaying”, due to the fact that such a function may decay slower at infinity than an $L^{2}(\mathbb{R}^{n})$ function does. On the other hand, even for compactly supported initial data, the results presented here show that, assuming additional integrability beyond that of an $L^{2}(\mathbb{R}^{n})$ function, one can obtain well-posedness statements for rougher initial data than one obtains from Strichartz estimates (for more on this see [27, Section 1.3]).
@@
The proof is almost completely analogous to that of [30, Theorem 1.4], although we rely on Theorems 8.1 and 5.1 instead of results about adapted Besov spaces. We briefly sketch the idea (see also [27, Section 6.2]).
@@
The main difference between Theorem 8.5 and the results in [30] concerns the second inclusion in (8.6) and (8.8). Although standard embeddings between Besov and Sobolev spaces, combined with the $\varepsilon$ loss in (8.5) and (8.7), imply that the spaces of initial data in (8.5) and (8.7) are not fundamentally larger than those in [30], or vice versa, $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$ satisfies sharp embeddings into the $L^{p}$-based Sobolev scale, cf. Theorem 6.3. By contrast, the adapted Besov spaces from [30] embed in a sharp manner into the standard Besov scale. Hence Theorems 8.5 and 6.3 show that $L^{p}$ regularity of the initial data is pointwise preserved in an optimal sense, whereas [30, Theorem 1.4] yields the corresponding statement for Besov regularity.
@@ Acknowledgments
Yung would like to thank Andreas Seeger for a helpful discussion that inspired the proof of Theorem 6.3.

### v3

@@ 6.3. Fractional integration
Although Theorem 6.4 only yields a strict improvement of (6.11) for $p\leq 2\leq r$, for general $p\leq r$ we can nonetheless improve a classical result about the mapping properties of Fourier integral operators. Recall the definition of Fourier integral operators in standard form, from Definition 2.12.
@@
(6.13) $T:\mathcal{L}_{W,s+m+\frac{n+1}{2}(\frac{1}{p}-\frac{1}{q})}^{q,p}(\mathbb{R}^{n})\rightarrow\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$
@@
    $2\leq p\leq q<\infty$ and $m=\frac{1}{q}-\frac{1}{p}-2s(q)$;
  3. (c)
    $1\leq p\leq 2\leq q<\infty$ and $m=n(\frac{1}{q}-\frac{1}{p})$.
Then
(6.14) $T:\mathcal{H}^{p}(\mathbb{R}^{n})\rightarrow L^{q}(\mathbb{R}^{n})$
is bounded.
@@
One obtains (6.13) by combining Theorem 6.4 with the boundedness of $T$ from $\mathcal{L}_{W,s+m}^{q,q}(\mathbb{R}^{n})=\mathcal{H}^{s+m,q}_{FIO}(\mathbb{R}^{n})$ to $\mathcal{L}_{W,s}^{q,q}(\mathbb{R}^{n})=\mathcal{H}^{s,q}_{FIO}(\mathbb{R}^{n})$. For $m=s=0$, this invariance is contained in [20, Theorem 6.10] in cases (2) and (3), and the techniques used there also allow one to treat operators as in (1). For general $m,s\in\mathbb{R}$, the mapping property $T:\mathcal{H}^{s+m,q}_{FIO}(\mathbb{R}^{n})\rightarrow\mathcal{H}^{s,q}_{FIO}(\mathbb{R}^{n})$ can be found in [29, Proposition 3.3 and Corollary 3.4] in cases (1) and (3). In case (2) it follows from [31, Proposition 3.3], after precomposing with the operator $\langle D\rangle^{-m}$.
In turn, (6.14) follows from (6.13), together with Sobolev embeddings. Indeed, in cases (a) and (b), Theorem 6.3 implies that
$\mathcal{H}^{p}(\mathbb{R}^{n})\subseteq\mathcal{L}_{W,-s(p)}^{q,p}(\mathbb{R}^{n})=\mathcal{L}_{W,s(q)+m+\frac{n+1}{2}(\frac{1}{p}-\frac{1}{q})}^{q,p}(\mathbb{R}^{n}).$
Hence (6.13) and the embedding $\mathcal{L}_{W,s(q)}^{q,p}(\mathbb{R}^{n})\subseteq L^{q}(\mathbb{R}^{n})$, again from Theorem 6.3, yield (6.14). On the other hand, for (c) one can combine Theorems 6.3 and 6.4 to obtain
$\mathcal{H}^{p}(\mathbb{R}^{n})\subseteq\mathcal{L}_{W,-s(p)}^{2,p}(\mathbb{R}^{n})\subseteq\mathcal{L}_{W,-n(\frac{1}{p}-\frac{1}{2})}^{2,2}(\mathbb{R}^{n})$
and
$\mathcal{L}_{W,-m-n(\frac{1}{p}-\frac{1}{2})}^{2,2}(\mathbb{R}^{n})\subseteq\mathcal{L}_{W,-m-n(\frac{1}{p}-\frac{1}{2})-\frac{n+1}{2}(\frac{1}{2}-\frac{1}{q})}^{2,q}(\mathbb{R}^{n})\subseteq L^{q}(\mathbb{R}^{n}).$
An application of (6.13) in the middle, which in fact boils down to $L^{2}$ theory, thus concludes the proof. ∎
@@
In the case of an operator as in (1) with symbol in $S^{m}_{1,0}$, the mapping property in (6.14) is well known (see e.g. [39, Section IX.6.15]), and from this one can derive the same statement for operators as in (3). Corollary 6.6 improves upon this result in several ways, since it allows for the larger class of $S^{m}_{1/2,1/2,1}$ symbols, removes the assumption that the symbol has compact spatial support, cf. (2), and yields stronger estimates through (6.13), due to the sharpness of the embeddings in Theorem 6.3. In particular, as follows from the proof of (6.14) in case (c) , Corollary 6.6 extends the improved fractional integration result mentioned in Remark 6.5, from the identity operator to general Fourier integral operators.
Remark 6.8.
Corollary 6.6 is sharp, in the sense that the exponent $q$ cannot be enlarged (or the parameter $m$ decreased) for general Fourier integral operators as in (1), (2) or (3). In fact, (6.13) implies (6.14), and the latter is already sharp, as is noted in [39, Section IX.6.16].
Corollary 6.6 also implies that the operator from Section 5.2, despite not being bounded on $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$, is bounded from $\mathcal{L}_{W,s+\frac{n+1}{2}(\frac{1}{p}-\frac{1}{q})}^{q,p}(\mathbb{R}^{n})$ to $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$ if $q>p$.
@@ 7.1. A discrete description of the $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$ norm
Throughout this section, for each $R\geq 2$, fix a maximal collection $V_{R}\subseteq S^{n-1}$ of unit vectors such that $|\nu-\nu^{\prime}|\geq R^{-1/2}$ for all $\nu,\nu^{\prime}\in V_{R}$. Then $V_{R}$ has approximately $R^{(n-1)/2}$ elements. Let $(\chi_{\nu})_{\nu\in V_{R}}\subseteq C^{\infty}(\mathbb{R}^{n}\setminus\{0\})$ be an associated partition of unity. More precisely, each $\chi_{\nu}$ is positively homogeneous of degree 0 and satisfies $0\leq\chi_{\nu}\leq 1$ and
@@
for all $\nu\in V_{R}$. Such a collection is straightforward to construct (see e.g. [39, Section IX.4]), and the collection $\{\mathcal{F}^{-1}(\chi_{\nu})\mid R\geq 2,\nu\in V_{R}\}$ is uniformly bounded in $L^{1}(\mathbb{R}^{n})$.
@@
The proof of (7.1) where $p=q$ is covered by [31, Proposition 4.1], at least for $R=2^{k}$ with $k\in{\mathbb{N}}$. For general $p$ and $q$, the statement follows from [32, Proposition 2.4], since $\|f\|_{L^{p}(\mathbb{R}^{n})}$ and the Besov space norm $\|f\|_{B^{0}_{p,q}(\mathbb{R}^{n})}$ are comparable, which in turn follows from the assumption that $f$ has frequency support in a dyadic annulus.
@@ 7.2. Decoupling for the sphere
For $\kappa\geq 1$ and $R\geq 2\kappa$, let $S_{R,\kappa}$ be a $\kappa R^{-1}$-neighborhood of the sphere $S^{n-1}$ in $\mathbb{R}^{n}$, and let $V_{R}$ be as in the previous subsection. Then $\theta_{\nu}:=\mathrm{supp}(\chi_{\nu})\cap S_{R,\kappa}$, for each $\nu\in V_{R}$, is a curved rectangle of dimensions approximately $R^{-1/2}\times\dots\times R^{-1/2}\times R^{-1}$ pointing in the direction of $\nu$. Of course, the exact size of the region depends on $\kappa$, but this parameter will be fixed below. Moreover, $S_{R,\kappa}=\cup_{\nu\in V_{R}}\theta_{\nu}$, and the $\theta_{\nu}$ have finite overlap.
@@ 8.1. Local smoothing
For $p=q$, the statement is contained in [31, Theorem 4.4]. For $q=2$, it is a consequence of [32, Theorem 1.1] and the standard embedding $W^{s+\alpha(p)+\varepsilon,p}(\mathbb{R}^{n})\subseteq B^{s+\alpha(p)+\varepsilon/2}_{p,p}(\mathbb{R}^{n})$. For general $q\geq 2$, it follows by combining the latter result with the embedding $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})\subseteq\mathcal{L}_{W,s}^{2,p}(\mathbb{R}^{n})$ from (6.1). ∎
@@
for an implicit constant independent of $f\in W^{s+\sigma(p)+\varepsilon,p}(\mathbb{R}^{n})$. Here $\sigma(p):=0$ for $2<p\leq 2n/(n-1)$, and $\sigma(p):=2s(p)-1/p$ for $p\geq 2n/(n-1)$. By the Sobolev embeddings for $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$ in Theorem 6.3 (1), (8.1) improves upon (8.2) for $p\geq 2(n+1)/(n-1)$ and $2\leq q\leq p$. Moreover, by the sharpness of the Sobolev embedding in Theorem 6.3 (2), (8.1) in fact yields a strict improvement of (8.2) for such $p$ and $q$. On the other hand, for 2<p<2(n+1)/(n1)2<p<2(n+1)/(n-1) and $2\leq q\leq p$, (8.1) neither follows from (8.2), nor vice versa. In particular, due to the sharpness of the Sobolev embedding in Theorem 6.3 (2), (8.1) yields sharper estimates than (8.2) for certain initial data. Here it is relevant to note that the exponents $\alpha(p)$ in (8.1) and $\sigma(p)$ in (8.2) are sharp, for all $2<p<\infty$ and $2\leq q\leq p$. In the case of (8.1), this follows from (6.1) and [31, Theorem 5.3].
@@
By Theorem 6.4, one has $W^{1/2,2}(\mathbb{R}^{n})\subsetneq\mathcal{L}_{W,0}^{2,p}(\mathbb{R}^{n})$ for p=2(n+1)/(n1)p=2(n+1)/(n-1). Given that $\alpha(p)=0$, Theorem 8.1 therefore almost yields a strict improvement of the classical Strichartz estimate
@@ 8.2. Nonlinear wave equations
In this subsection we indicate how our results can be applied to nonlinear wave equations with initial data in $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$ for $p>2$. Such initial data might be referred to as “slowly decaying”, due to the fact that such a function may decay slower at infinity than an $L^{2}(\mathbb{R}^{n})$ function does. On the other hand, even for compactly supported initial data, the results presented here show that, assuming additional integrability beyond that of an $L^{2}(\mathbb{R}^{n})$ function, one can obtain well-posedness statements for rougher initial data than one obtains from Strichartz estimates (for more on this see [29, Section 1.3]).
@@
The proof is almost completely analogous to that of [32, Theorem 1.4], although we rely on Theorems 8.1 and 5.1 instead of results about adapted Besov spaces. We briefly sketch the idea (see also [29, Section 6.2]).
@@
The main difference between Theorem 8.5 and the results in [32] concerns the second inclusion in (8.6) and (8.8). Although standard embeddings between Besov and Sobolev spaces, combined with the $\varepsilon$ loss in (8.5) and (8.7), imply that the spaces of initial data in (8.5) and (8.7) are not fundamentally larger than those in [32], or vice versa, $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$ satisfies sharp embeddings into the $L^{p}$-based Sobolev scale, cf. Theorem 6.3. By contrast, the adapted Besov spaces from [32] embed in a sharp manner into the standard Besov scale. Hence Theorems 8.5 and 6.3 show that $L^{p}$ regularity of the initial data is pointwise preserved in an optimal sense, whereas [32, Theorem 1.4] yields the corresponding statement for Besov regularity.
@@ Acknowledgments
The authors would like to thank the referee for carefully reading the manuscript, and for several useful remarks and suggestions. Rozendaal would like to thank Yiyu Tang for pointing out various typos and minor mistakes. Yung would like to thank Andreas Seeger for a helpful discussion that inspired the proof of Theorem 6.3. Yung is partially supported by a Future Fellowship FT200100399 from the Australian Research Council.